Use the identities to find the exact value of the remaining trigonometric functions if sin x = -5/13 and tan x = -5/12.
Answer by chibisan(131)

sinx = opp/hyp
tanx = opp/adj
since all the sides are given, the remaining trigonometic function ;
cosx = adj/hyp
cosx= 12/13 {ans}